# Event Camera Deraining

### v1

@@ Abstract
Event cameras excel in high temporal resolution and dynamic range but suffer from dense noise in rainy conditions. Existing event deraining methods face trade-offs between temporal precision, deraining effectiveness, and computational efficiency. In this paper, we propose PRE-Mamba, a novel point-based event camera deraining framework that fully exploits the spatiotemporal characteristics of raw event and rain. Our framework introduces a 4D event cloud representation that integrates dual temporal scales to preserve high temporal precision, a Spatio-Temporal Decoupling and Fusion module (STDF) that enhances deraining capability by enabling shallow decoupling and interaction of temporal and spatial information, and a Multi-Scale State Space Model (MS3M) that captures deeper rain dynamics across dual-temporal and multi-spatial scales with linear computational complexity. Enhanced by frequency-domain regularization, PRE-Mamba achieves superior performance (0.95 SR, 0.91 NR, and 0.4s/M events) with only 0.26M parameters on EventRain-27K, a comprehensive dataset with labeled synthetic and real-world sequences. Moreover, our method generalizes well across varying rain intensities, viewpoints, and even snowy conditions.
@@ 1 Introduction
Event cameras have revolutionized mobile applications with their high temporal resolution (microsecond level), high dynamic range (>120 dB), and low power consumption (<10 mW), driving breakthroughs in high-speed tracking [1, 2, 3], SLAM [4, 5, 6], and obstacle avoidance [7, 8]. Unlike conventional frame cameras, they asynchronously capture pixel-level intensity changes without global exposure, eliminating motion blur even at extreme speeds. However, despite excelling in dynamic scenes, event cameras suffer significant performance degradation under adverse weather, particularly in rain [9, 10]. The rapid motion of rain streaks triggers excessive intensity changes, generating dense noise that overwhelms valid scene information and exacerbates sampling rates. This severely impacts the accuracy and latency of downstream tasks, limiting real-world reliability.
@@
To address this, we propose PRE-Mamba, the first Point-based deRaining framework for Event camera. Our framework first introduces a 4D event cloud representation (§3.1), integrating inter- and intra-temporal windows to preserve high temporal precision. To handle the spatial irregularity and motion characteristics of rain, we propose a Spatio-Temporal Decoupling and Fusion module (STDF, §3.2), enabling shallow decoupling and interaction of temporal and spatial information. For deeper modeling, we propose a Multi-Scale State Space Model (MS3M, §3.3) that efficiently captures complex rain dynamics across dual temporal and multi-spatial scales with linear computational complexity. To improve feature discrimination, we introduce a frequency regularization term (§3.4) in the loss function, guiding the model to learn rain distribution patterns in the frequency domain. To evaluate our approach, we construct EventRain-27K  (§4), the first comprehensive point-based event deraining dataset, comprising 18K  labeled synthetic and 9K unlabeled real-world rain sequences. Our model achieves state-of-the-art performance (0.95 SR, 0.91 NR, and 0.4s/M events) with only 0.26M parameters, validated through extensive experiments (§5). Moreover, it generalizes well across diverse rain intensities and snowy conditions. Our contributions are summarized as follows:
@@ 2 Related Work
Existing deraining approaches for event camera. Research on event-based deraining remains limited despite its importance for outdoor applications. Existing methods, such as EGVD [32], RainVID&SS [33], and NESD  [34], leverage the high temporal resolution and dynamic range of event cameras to assist frame camera by providing prior knowledge of rain location and density. Other methods designed for event-based deraining typically convert asynchronous event streams into synchronous frame-like representations, aligning with image deraining but sacrificing temporal resolution and sparsity. For instance, DistillNet [10] partially preserves temporal resolution through voxel representations but still loses their asynchronous nature. Moreover, the unique motion and spatiotemporal distribution of rain create a domain barrier, limiting the applicability of traditional event denoising methods like filtering [35, 36, 37] and clustering [38]. In addition, existing deraining datasets primarily focus on frame camera [39, 40, 41, 42] or represent event data in image-like formats [43], further hinder the event camera deraining research.
@@
Event Data Processing Architectures. Event data processing approaches are broadly categorized into frame-based and point-based frameworks. Frame-based methods [44, 45, 37, 46] not only face the aforementioned limitations but also incur redundant computational overhead by processing event-free regions [47], making them inefficient for resource-constrained environments [45, 48]. In contrast, point-based methods process events in their native form, preserving temporal resolution and sparsity. These include Point Networks [49, 50], Graph Neural Networks (GNNs) [51], and Spiking Neural Networks (SNNs) [21, 52], excel in capturing rapid motion and fine-grained temporal dynamics. However, GNNs and SNNs often require specialized hardware or face performance limitations, while traditional Point Cloud networks, designed with assumptions of permutation and transformation invariance, are ill-suited for temporal event data, necessitating specialized feature extraction for time-aware processing.
@@ 3 Method
To preserve fine-grained spatiotemporal information while enabling efficient processing, we transform raw event streams into a structured 4D event cloud representation. The raw event data is represented as a sequence =$\mathcal{E}=\{e_{i}\}_{i=1}^{N}$, where each event $e_{i}=(x_{i},y_{i},t_{i},p_{i})$ consists of spatial coordinates $(x_{i},y_{i})$, timestamp $t_{i}$, and polarity $p_{i}$.
@@ 3.2 Spatio-Temporal Decoupling and Fusion
Finally, the temporally-modulated features are combined with the original spatial and temporal representations through residual addition ($\mathcal{R}$). The integrated features are then processed sequentially through 1D convolution ($\Phi_{1d}$), batch normalization (BN), and activation functions ($\sigma$) to ensure training stability:
@@ 3.3 Multi-Scale State Space Model
Dual-Temporal-Scale Global Modeling: While Mamba excels in sequence modeling, it struggles to leverage temporal features [61] for event deraining. Inspired by the Motion-aware State Space Model (MSSM), which excels in LiDAR-based moving object segmentation [62], we integrate MSSM into our dual-temporal-scale architecture. This enables explicit modeling of rain events’ appearance and motion, improving rain-background separation.
The architecture mainly consists of two complementary branches: an intra-window branch processes spatial appearance features $f_{\text{s}}^{\text{intra}}$ through Reversed Aggregation (RA) and 1D convolution, capturing rain streak patterns and structural details for rain-background separation; an inter-window branch extracts motion-aware features $f_{\text{m}}^{\text{inter}}$ by analyzing temporal dependencies, enabling the model to separate dynamic rain streaks from moving objects or scene changes. An adaptive gating mechanism [63] generates weight factors $f_{\text{G}}$ to emphasize rain-relevant features. Features from both branches are fused via cross-product attention, combined with inter-branch interactions, and fed into the SSM to learn global features through scanning [64], followed by gating signal modulation:
@@ 3.4 Training Loss
where $\mathcal{F}(\cdot)$ denotes the 1D FFT, $L$ is the sequence length, and $\epsilon$, $\epsilon^{\prime}$ ensure numerical stability, and $Y$ and $P$ denote the ground truth labels and predicted outputs, respectively. Unlike traditional $\mathcal{L}_{\text{fft}}$ used in image reconstruction [67, 53], our formulation operates directly on event labels, leveraging event cameras’ microsecond-level resolution to capture rain dynamics. Rainfall intensity directly correlated to event density and label $1$’s density (rain events labeled as $1$). As illustrated in Figure 3, high-density rainfall manifests as low-frequency continuous patterns in the label sequence, while low-density rain exhibits high-frequency sparse patterns. This spectral distinction enables $\mathcal{L}_{\text{fft}}$ to align predictions with physically consistent global patterns.
@@ 5.1 Experiments Setups
[…]
@@ 5.2 Comparisons with State-of-The-Art Methods
DWF[74] N/A N/A 0.0954 1.36$\times$
KNoise[75] N/A N/A 0.0198 6.55$\times$
YNoise[76] N/A N/A 0.0513 2.53$\times$
RED[77] N/A N/A 2.2716 0.06$\times$
EDnCNN[77] 234.51 614.55K 20.1885 1.0$\times$
AEDNet[78] 4400.46 45.87M 43.4250 0.46$\times$
DistillNet[10] 255.17 18.96M 0.2029 99.50$\times$
EDformer[58] 8.41 49.80K 2.4943 8.09$\times$
Ours 6.23 264.63K 0.0987 204.54$\times$
@@
Model Complexity and Efficiency Comparison. In Table 2, we compare the computational efficiency of our method and baselines in terms of FLOPs, parameters, and inference time for processing 100K events on an NVIDIA RTX A6000 GPU. Filtering-based methods enable fast inference but suffer from poor deraining performance due to lack of task-specific optimization. EDnCNN [77] and AEDNet [78] incur high computational overhead from encoding or sampling spatial-temporal neighborhoods, resulting in high FLOPs and slow inference. Transformer-based EDformer [58] reduces FLOPs and parameters but suffers from high latency due to the quadratic complexity of its attention mechanism. DistillNet [10] achieves a lower inference time by voxelizing events but at the cost of high computational complexity, large model size, and loss of fine-grained temporal precision. In contrast, our method achieves SOTA performance with only 6.23G FLOPs and 0.26M parameters, processing 100K events in 0.0987s and scaling to 1M events in 0.398s. Compared to EDnCNN and AEDNet, our method reduces FLOPs to 2.66% and 0.14% of theirs, respectively, while achieving 204.5$\times$ and 440.0$\times$ faster inference. Compared to the lightweight EDformer, our method achieves a 25.3$\times$ speedup with comparable model size, excelling in both accuracy and computational efficiency.
@@ 5.4 Generalization to Snowy Datasets
While our primary focus is on event-based deraining, we further evaluate our method’s generalization capability on snow removal tasks. As shown in Figure 7 (last two rows), our approach achieves promising performance in snow removal without any architectural modifications, highlighting its inherent ability to handle diverse weather patterns. More visual comparisons are provided in the Appendix.
@@ 6 Conclusion
To the best of our knowledge, PRE-Mamba is the first point-based framework designed for event camera deraining. It features a 4D event cloud representation that preserves native temporal resolution through inter- and intra-temporal windows, a spatiotemporal decoupling and fusion module for effective feature extraction, and a multi-scale state space model that captures complex rain dynamics across dual temporal and multi-spatial scales with linear complexity. To enhance learning, we introduce a frequency regularization term combined with cross-entropy loss, guiding the model to learn rain distribution patterns in the frequency domain. Furthermore, we contribute a large-scale event deraining dataset with diverse rain intensities and scenarios. Extensive experiments show SOTA performance on synthetic and real-world datasets, with high efficiency and a lightweight design for real-time applications.

### v2

@@ Abstract
Event cameras excel in high temporal resolution and dynamic range but suffer from dense noise in rainy conditions. Existing event deraining methods face trade-offs between temporal precision, deraining effectiveness, and computational efficiency. In this paper, we propose PRE-Mamba, a novel point-based event camera deraining framework that fully exploits the spatiotemporal characteristics of raw event and rain. Our framework introduces a 4D event cloud representation that integrates dual temporal scales to preserve high temporal precision, a Spatio-Temporal Decoupling and Fusion module (STDF) that enhances deraining capability by enabling shallow decoupling and interaction of temporal and spatial information, and a Multi-Scale State Space Model (MS3M) that captures deeper rain dynamics across dual-temporal and multi-spatial scales with linear computational complexity. Enhanced by frequency-domain regularization, PRE-Mamba achieves superior performance (0.95 SR, 0.91 NR, and 0.4s/M events) with only 0.26M parameters on EventRain-27K, a comprehensive dataset with labeled synthetic and real-world sequences. Moreover, our method generalizes well across varying rain intensities, viewpoints, and even snowy conditions. Code and dataset: https://github.com/softword-tt/PRE-Mamba.
@@ 1 Introduction
Event cameras have revolutionized mobile applications with their high temporal resolution (microsecond level), high dynamic range (>120 dB), and low power consumption (<10 mW) [1, 2], driving breakthroughs in high-speed tracking [3, 4], SLAM [5, 6, 7], localization [8] and obstacle avoidance [9, 10]. Unlike frame cameras, they asynchronously capture pixel-level intensity changes without global exposure, eliminating motion blur even at extreme speeds. However, despite excelling in dynamic scenes, event cameras suffer significant performance degradation under adverse weather, particularly in rain [11, 12]. The rapid motion of rain streaks triggers excessive intensity changes, generating dense noise that overwhelms valid scene information and exacerbates sampling rates. This severely impacts accuracy and latency in downstream tasks on drones [13, 14, 15] and autonomous vehicles [16, 17, 18], limiting real-world reliability.
@@
To address this, we propose PRE-Mamba, the first Point-based deRaining framework for Event camera. As a point-based method, it processes raw events directly instead of quantized frames, fully preserving the sensor’s native microsecond (µs) temporal resolution and facilitating downstream high-speed vision tasks. Our framework first introduces a 4D event cloud representation (§3.1), integrating inter- and intra-temporal windows to preserve high temporal precision. To handle the spatial irregularity and motion characteristics of rain, we propose a Spatio-Temporal Decoupling and Fusion module (STDF, §3.2), enabling shallow decoupling and interaction of temporal and spatial information. For deeper modeling, we propose a Multi-Scale State Space Model (MS3M, §3.3) that efficiently captures complex rain dynamics across dual temporal and multi-spatial scales with linear computational complexity. To improve feature discrimination, we introduce a frequency regularization term (§3.4) in the loss function, guiding the model to learn rain distribution patterns in the frequency domain. To evaluate our approach, we construct EventRain-27K  (§4), the first comprehensive point-based event deraining dataset, comprising 18K  labeled synthetic and 9K unlabeled real-world rain sequences. Our model achieves state-of-the-art performance (0.95 SR, 0.91 NR, and 0.4s/M events) with only 0.26M parameters, validated through extensive experiments (§5). Moreover, it generalizes well across diverse rain intensities and snowy conditions. Our contributions are summarized as follows:
@@ 2 Related Work
Existing deraining approaches involving event camera. Research on event-based deraining remains limited despite its importance for outdoor applications. Existing methods, such as EGVD [40], RainVID&SS [41], and NESD  [42], leverage the high temporal resolution and dynamic range of event cameras to assist frame camera by providing prior knowledge of rain location and density. Other methods designed for event-based deraining typically convert asynchronous event streams into synchronous frame-like representations, aligning with image deraining but sacrificing temporal resolution and sparsity. For instance, DistillNet [11] partially preserves temporal resolution through voxel representations but still loses their asynchronous nature. Moreover, the unique motion and spatiotemporal distribution of rain create a domain barrier, limiting the applicability of traditional event denoising methods like filtering [43, 44, 45] and clustering [46]. In addition, existing deraining datasets primarily focus on frame camera [47, 48, 49, 50] or represent event data in image-like formats [51], further hinder the event camera deraining research.
@@
Event Data Processing Architectures. Event data processing approaches are broadly categorized into frame-based and point-based frameworks. Frame-based methods [52, 53, 45, 54] not only face the aforementioned limitations but also incur redundant computational overhead by processing event-free regions [55], making them inefficient for resource-constrained environments [53, 56]. In contrast, point-based methods process events in their native form, preserving temporal resolution and sparsity. These include Point Networks [57, 58], Graph Neural Networks (GNNs) [59], and Spiking Neural Networks (SNNs) [29, 60], excel in capturing rapid motion and fine-grained temporal dynamics. However, GNNs and SNNs often require specialized hardware or face performance issues, while Point Cloud networks, based on assumptions of permutation and transformation invariance, are poorly suited for temporal event data and require time-aware feature extraction.
@@ 3 Method
To preserve fine-grained spatiotemporal information while enabling efficient processing, we transform raw event streams into a structured 4D event cloud representation. The raw event data is represented as a sequence $\mathcal{E}=\{e_{i}\}_{i=1}^{N}$, where each event $e_{i}=(x_{i},y_{i},t_{i},p_{i})$ consists of spatial coordinates $(x_{i},y_{i})$, timestamp $t_{i}$, and polarity $p_{i}$.
@@ 3.2 Spatio-Temporal Decoupling and Fusion
Finally, the temporally-modulated features are combined with the original spatial and temporal representations via residual addition ($\mathcal{R}$). The integrated features are sequentially processed by 1D convolution ($\Phi_{1d}$), batch normalization (BN), and activation ($\sigma$) to ensure training stability:
@@ 3.3 Multi-Scale State Space Model
Dual-Temporal-Scale Global Modeling: While Mamba excels in sequence modeling, it struggles to leverage temporal features [69] for event deraining. Inspired by the Motion-aware State Space Model (MSSM), which excels in LiDAR-based moving object segmentation [70], we integrate MSSM into our dual-temporal-scale architecture. This enables explicit modeling of rain events’ appearance and motion, improving rain–background separation.
The architecture mainly consists of two complementary branches: an intra-window branch processes spatial appearance features $f_{\text{s}}^{\text{intra}}$ through Reversed Aggregation (RA) and 1D convolution, capturing rain patterns and structural details for rain-background separation; an inter-window branch extracts motion-aware features $f_{\text{m}}^{\text{inter}}$ by analyzing temporal dependencies, enabling separation of dynamic rain streaks from moving objects or scene changes. An adaptive gating mechanism [71] generates weight $f_{\text{G}}$ to emphasize rain-relevant features. Outputs from both branches are fused via cross-product attention with inter-branch interaction, then passed to SSM for global feature learning via scanning [72], followed by gating signal modulation.
@@ 3.4 Training Loss
where $\mathcal{F}(\cdot)$ denotes the 1D FFT, $L$ is the sequence length, and $\epsilon$, $\epsilon^{\prime}$ ensure numerical stability, and $Y$ and $P$ denote the ground truth labels and predicted outputs, respectively. Unlike traditional $\mathcal{L}_{\text{fft}}$ used in image reconstruction [75, 61], our formulation operates directly on event labels, leveraging event cameras’ microsecond-level resolution to capture rain dynamics. Rainfall intensity directly correlated to event density and label $1$’s density (rain events labeled as $1$). As illustrated in Figure 3, high-density rainfall manifests as low-frequency continuous patterns in the label sequence, while low-density rain exhibits high-frequency sparse patterns. This spectral distinction enables $\mathcal{L}_{\text{fft}}$ to align predictions with physically consistent global patterns [76].
@@ 5.1 Experiments Setups
Evaluation Metrics. We use event Denoising Accuracy (DA), Signal Retention (SR) and Noise Removal (NR) to measure the event rain removal performance on PRE-Mamba. There is $\text{DA}=\frac{1}{2}(\text{SR}+\text{NR})=(\frac{\text{PB}}{\text{TB}}+\frac{\text{PR}}{\text{TR}})$, where $\text{PB},\text{TB},\text{PR},\text{TR}$ are the predicted background, the predicted rain, the real background, and the number of real rain, respectively. DA measures rain removal while preserving real motion (e.g., vehicles, pedestrians) and structural details (e.g., roads, building profiles). SR and NR analyze signal retention and noise suppression, respectively.
@@ 5.2 Comparisons with State-of-The-Art Methods
[…]
@@
[…]
@@ 5.4 Generalization to Snowy Datasets
While primarily designed for event-based deraining, our method also generalizes well to snow removal, as demonstrated in Figure 7 (last two rows). Without architectural changes, it delivers promising results under snowy conditions, showcasing robustness to diverse weather patterns. Additional visual results are included in the Appendix.
@@ 6 Conclusion
To the best of our knowledge, PRE-Mamba is the first point-based framework for event camera deraining. It adopts a 4D event cloud to preserve native temporal resolution via inter- and intra-temporal windows, a spatiotemporal decoupling and fusion module for efficient feature extraction, and a multi-scale state space model that captures rain dynamics across spatial and temporal scales with linear complexity. A frequency regularization term and cross-entropy loss guide learning, and we contribute a large-scale dataset with diverse rain scenarios. Extensive experiments demonstrate SOTA performance on synthetic and real-world data. The method is efficient, lightweight, and real-time capable, with potential for robust multi-agent perception [88, 89].
Acknowledgments This paper was supported by the Natural Science Foundation of China under Grant 62371269, Guangdong Innovative and Entrepreneurial Research Team Program (2021ZT09L197), and Meituan Academy of Robotics Shenzhen.